# Slicing Wasserstein over Wasserstein via Functional Optimal Transport

### v1

@@ 1 Introduction
Optimal transport (OT) enables geometrically meaningful Wasserstein distances between probability measures on arbitrary Polish spaces $\mathcal{X}$ (Villani, 2003), while remaining computationally tractable. This has had a profound impact on machine learning, where Wasserstein distances are used to train neural networks (Tong et al., 2024) and to compare data distributions (Yang et al., 2019). A key feature of OT is its applicability to non-Euclidean spaces, even allowing the definition of Wasserstein distances on Wasserstein spaces $\mathcal{P}_{2}(\mathcal{X})$ (Bonet et al., 2025b). This is particularly useful for comparing distributions over non-Euclidean objects. For example, Euclidean distances as ground metric between two images often yield poor results (Stanczuk et al., 2021), whereas Wasserstein distances are robust to small image perturbations (Beckmann et al., 2025). Similarly, comparing point clouds (Nguyen et al., 2021) is natural with OT but not even well-defined with Euclidean distances. While most OT applications focus either on comparing pairwise objects or distributions over Euclidean spaces, recent work leverages Wasserstein distances on Wasserstein spaces for non-Euclidean domains, such as image (Dukler et al., 2019) or point cloud (Haviv et al., 2025) distributions.
@@
The underlying concept of multilevel OT has been introduced under various names, including hierarchical OT (Schmitzer & Schnörr, 2013; Lee et al., 2019), mixture Wasserstein (Chen et al., 2018; 2019; Delon & Desolneux, 2020), and Wasserstein over Wasserstein (WoW) (Bonet et al., 2025b). It has applications beyond images and shapes, including domain adaptation (Lee et al., 2019; El Hamri et al., 2022), single-cell analysis (Lin et al., 2023), point cloud registration (Steuernagel et al., 2023), Bayesian inference (Nguyen & Mueller, 2024), generative modelling (Atanackovic et al., 2025; Haviv et al., 2025), document analysis (Yurochkin et al., 2019), Gromov–Wasserstein approximations (Mémoli, 2011; Piening & Beinert, 2025a), and reinforcement learning (Ziesche & Rozo, 2023). Extending this framework with another Polish space $\mathcal{Y}$ for dataset labels yields the OT dataset distance (OTDD), defined on $\mathcal{P}_{2}(\mathcal{Y}\times\mathcal{P}_{2}(\mathcal{X}))$ (Alvarez-Melis & Fusi, 2020). However, all these approaches incur high computational cost due to repeated pairwise Wasserstein evaluations.
@@
Due to the complexity of OT, sliced Wasserstein distances (Bonneel et al., 2015) provide efficient OT-based alternatives to standard Wasserstein distances. Initially developed for probability measures on Euclidean spaces, they have since been extended to the sphere (Bonet et al., 2023a; Quellmalz et al., 2023), manifolds (Bonet et al., 2025a), functions (Garrett et al., 2024), hyperbolic spaces (Bonet et al., 2023b), the rotation group (Quellmalz et al., 2024), and matrices (Bonet et al., 2023c). For WoW-type distances, sliced accelerations have been proposed for Gaussian mixtures (Nguyen & Mueller, 2024; Piening & Beinert, 2025b) and more generally for measures over measures (meta-measures) via the sliced OTDD (s-OTDD) (Nguyen et al., 2025). The s-OTDD employs a hierarchical slicing approach based on the method of moments. However, it is only well-defined for finite moments, and practical implementations are limited to the first few moments because of numerical instability – originally, the first five.
@@
  • Beyond 1d meta-measures, we extend our approach to the multivariate case by introducing the double-sliced Wasserstein (DSW) metric between meta-measures. Illustrating the usefulness of our DSW metric as a WoW replacement, we prove a form of topological metric equivalence between the two for discretized meta-measures.
@@ 2 Wasserstein Distances
The so-called Wasserstein distance or Kantorovich–Rubinstein metric is an optimal transport-based similarity gauge between probabilities on a common Polish space. To this end, let $\mathcal{X}$ be a Polish space, let $\mathcal{P}(\mathcal{X})$ be the space of Borel probability measures on $\mathcal{X}$ with respect to the Borel $\sigma$-algebra induced by the underlying metric, and let $\mathcal{P}_{p}(\mathcal{X})$, $p\in[1,\infty)$, be the subset of measures with finite $p$th moment. For $\mu\in\mathcal{P}(\mathcal{X})$ and a second Polish space $\mathcal{Y}$, the push-forward by a measurable map $T\colon\mathcal{X}\to\mathcal{Y}$ is defined by $T_{\sharp}\,\mu\coloneqq\mu\circ T^{{-}1}\in\mathcal{P}(\mathcal{Y})$. The set of transport plans between $\mu\in\mathcal{P}(\mathcal{X}),\nu\in\mathcal{P}(\mathcal{Y})$ is given by
@@
defines a metric on $\mathcal{P}_{2}(\mathcal{X})$. More precisely, $(\mathcal{P}_{2}(\mathcal{X}),\operatorname{W})$ is again a complete separable metric space, allowing the construction of the so-called Wasserstein over Wasserstein (WoW) distance ${\operatorname{\mathbf{W}}}(\cdot,\cdot;\mathcal{X})\coloneqq\operatorname{W}(\cdot,\cdot;\mathcal{P}_{2}(\mathcal{X}))$, which is studied in (Bonet et al., 2025b).
In difference to other similarity gauges like the Kullback–Leibler divergence or the total variation, Wasserstein distances leverage the underlying geometry, allowing for meaningful comparisons between empirical measures. Although the Wasserstein distance relies on a linear program, the actual calculation is computationally costly. For two empirical measures supported at $n$ points in $\mathbb{R}^{d}$ equipped with the Euclidean metric, the exact computation has complexity $\mathcal{O}(n^{3}\log n)$. The approximate computation based on entropic regularization still has complexity $\mathcal{O}(n^{2}\log n)$, see (Peyré & Cuturi, 2019). Notably, this computational burden becomes even more involved for non-Euclidean metric spaces, where the computation of the underlying distance itself is challenging. For instance, the computation of the WoW distance relies on the pointwise evaluations of Wasserstein distances. If the empirical meta-measures in $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ are supported on $N$ empirical measures on $\mathcal{P}_{2}(\mathbb{R}^{d})$, each with $n$ support points, then the approximate calculation of the required distance matrix already has complexity $\mathcal{O}(N^{2}n^{2}\log n)$.
@@ 3.1 Slicing Infinite Dimensional Banach Spaces
As preliminary step towards an implementable sliced WoW distance, we consider the slicing on an infinite dimensional, separable Banach space $U$ with dual $U^{*}$. Relying on the dual pairing, we generalize the slicing operator equation 5 by
$\pi_{v}\colon U\to\mathbb{R},\;u\mapsto\langle v,u\rangle,\qquad v\in U^{*}.$ (7)
The crucial point in defining a sliced distance on $\mathcal{P}_{2}(U)$ is that there exists no uniform probability measure on the infinite dimensional sphere. As remedy, we choose an arbitrary $\xi\in\mathcal{P}_{2}(U^{*})$ and define the $\xi$-based sliced Wasserstein distance as
@@
This approach extends the slicing on Hilbert spaces (Han, 2023). However, unlike (Han, 2023), we do not construct specific measures on the sphere. This allows the use of easy-to-sample slicing measures. If the support of $\xi$ covers all directions in $U^{*}$, the $\xi$-based SW distance becomes a metric. The proof is given in Appendix A.
@@
For $\xi\in\mathcal{P}_{2}(U^{*})$, the $\xi$-based SW distance defines a well-defined pseudo-metric. If $\operatorname{supp}\xi\cap\operatorname{span}v\not\in\bigl\{\emptyset,\{0\}\bigr\}$ for all $v\in U^{*}$, then equation 8 defines a metric on $\mathcal{P}_{2}(U)$.
If the slicing measure $\xi$ has full support, then the assumption in Theorem 3.1 is fulfilled, and the $\xi$-based SW distance is a metric. Two measures $\xi_{1},\xi_{2}\in\mathcal{P}_{2}(U^{*})$ are equivalent if they are mutually absolutely continuous. If their Radon–Nikodým derivatives ${\,\mathrm{d}\xi_{1}}/{\,\mathrm{d}\xi_{2}}$ and ${\,\mathrm{d}\xi_{2}}/{\,\mathrm{d}\xi_{1}}$ are bounded, then the resulting SW distances are metrically equivalent. The proof is again given in Appendix A.
@@
In the finite-dimensional Euclidean setting, special cases of the $\xi$-based SW distance correspond, for instance, to so-called energy measures on the sphere (Nguyen & Ho, 2023) and the standard Gaussian (Nadjahi et al., 2021). Relying on the latter, we obtain a strong equivalence to original SW equation 6 if $\xi$ is equivalent to the standard Gaussian. The short proof is given in Appendix A.
@@
$c_{1}\operatorname{SW}(\mu,\nu;\xi)\leq\operatorname{SW}(\mu,\nu)\leq c_{2}\operatorname{SW}(\mu,\nu;\xi)\quad\forall\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}).$ (10)
@@ 3.2 Slicing the 1D Wasserstein Space
For the later implementation, we require an easy-to-sample slicing measure. To this end, we propose to use Gaussians, i.e., measures $\xi\in\mathcal{P}_{2}(L_{2}([0,1]))$ such that $\pi_{v,\sharp}\xi$ is Gaussian for all $v\in L_{2}([0,1])$, see (Bogachev, 1998). On $L_{2}([0,1])$, there exists a one-to-one correspondence between Gaussian measures and Gaussian processes (Rajput & Cambanis, 1972, Thm. 2). In our numerics, we restrict ourselves to the Gaussian process $G$ that is related to the covariance kernel
@@
Since the kernel is smooth, the sample paths (realizations) of $G$ are smooth too, i.e., $G~\in~\mathcal{C}^{\infty}([0,1])$ almost surely, see (Costa et al., 2023, Cor. 1). Since $k_{\sigma}$ is universal (Steinwart, 2001), the corresponding Gaussian measure has full support (Van Der Vaart et al., 2008).
@@ 4 Double-Slicing the Wasserstein Space
where we again integrate with respect to the uniform measure. Based on finite-dimensional slicing, $\operatorname{\mathbf{SW}}$} essentially reduces meta-measures $\bm{\mu}$ and $\bm{\nu}$ to a series of 1d meta-measures $\pi_{\theta,\sharp}\bm{\mu}$ and $\pi_{\theta,\sharp}\bm{\nu}$.
As the computation of the 1d WoW distance remains challenging, we resort to a hierarchical slicing approach, inspired by (Nguyen et al., 2025; Piening & Beinert, 2025b). Using quantile slicing equation 11 with slicing measure $\xi\in\mathcal{P}_{2}(L_{2}([0,1]))$, we introduce the double-sliced WoW distance
@@
Numerically, $\operatorname{\mathbf{DSW}}$} can be implemented combining several integration techniques. In the following, we consider the Gaussian $\xi$ related to the kernel in equation 12 and empirical meta-measures $\bm{\mu},\bm{\nu}\in\mathcal{P}_{\mathrm{e}}(\mathcal{P}_{\mathrm{e}}(\mathbb{R}^{d}))$ as in equation 17. To approximate the outer integral over $\mathbb{S}^{d-1}$ and the inner integral over $\xi$ simultaneously, we employ the Monte Carlo method. For a sample path $g$ of the corresponding Gaussian process and a direction $\theta\in\mathbb{S}^{d-1}$, we evaluate the integrand as follows: The first slicing gives $\bm{\pi}_{\theta,\sharp}\bm{\mu}=(1/N)\sum{}_{i=1}^{N}\delta_{\pi_{\theta,\sharp}\mu_{i}}$, and the quantile mapping $q_{\sharp}\bm{\pi}_{\theta,\sharp}\bm{\mu}=(1/N)\sum{}_{i=1}^{N}\delta_{q(\pi_{\theta,\sharp}\mu_{i})}$, where the piecewise constant quantile functions can be determined by sorting the support points. For the slicing on $L_{2}([0,1])$, we employ a quadrature to approximate the inner product. To this end, for knots $t_{1},\dots,t_{R}\in[0,1]$ and weights $w_{1},\dots,w_{R}$, we estimate
$\widehat{\pi_{g,\sharp}q_{\sharp}\bm{\pi}_{\theta,\sharp}\mu}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\widehat{\langle q(\pi_{\theta,\sharp}\mu_{i}),g\rangle}}\quad\text{with}\quad\widehat{\langle q(\pi_{\theta,\sharp}\mu_{i}),g\rangle}=\sum_{r=1}^{R}w_{r}\,q(\pi_{\theta,\sharp}\mu_{i})(t_{r})\,g(t_{r}).$ (19)
@@
$\widehat{\text{DSW}}(\bm{\mu},\bm{\nu})\coloneqq\biggl({\frac{1}{S}\sum_{s=1}^{S}\operatorname{W}^{2}\bigl(\widehat{\pi_{g_{s},\sharp}q_{\sharp}\bm{\pi}_{\theta_{s},\sharp}\mu},\widehat{\pi_{g_{s},\sharp}q_{\sharp}\bm{\pi}_{\theta_{s},\sharp}\nu};\mathbb{R}\bigr)}\biggr)^{\frac{1}{2}}.$ (20)

### v2

@@ 1 Introduction
Optimal transport (OT) enables geometrically meaningful Wasserstein distances between probability measures on arbitrary Polish spaces $\mathcal{X}$ (Villani, 2003), while remaining computationally tractable. This has had a profound impact on machine learning, where Wasserstein distances are used to train neural networks (Tong et al., 2024) and to compare data distributions (Yang et al., 2019). A key feature of OT is its applicability to non-Euclidean spaces, even allowing the definition of Wasserstein distances on Wasserstein spaces $\mathcal{P}_{2}(\mathcal{X})$ (Bonet et al., 2025c). This is particularly useful for comparing distributions over non-Euclidean objects. For example, Euclidean distances as ground metric between two images often yield poor results (Stanczuk et al., 2021), whereas Wasserstein distances are robust to small image perturbations (Beckmann et al., 2025). Similarly, comparing point clouds (Nguyen et al., 2021b) is natural with OT but not even well-defined with Euclidean distances. While most OT applications focus either on comparing pairwise objects or distributions over Euclidean spaces, recent work leverages Wasserstein distances on Wasserstein spaces for non-Euclidean domains, such as image (Dukler et al., 2019) or point cloud (Haviv et al., 2025) distributions.
@@
The underlying concept of multilevel OT has been introduced under various names, including hierarchical OT (Schmitzer and Schnörr, 2013; Lee et al., 2019), mixture Wasserstein (Chen et al., 2018; 2019; Delon and Desolneux, 2020), and Wasserstein over Wasserstein (WoW) (Bonet et al., 2025c). It has applications beyond images and shapes, including domain adaptation (Lee et al., 2019; El Hamri et al., 2022), single-cell analysis (Lin et al., 2023), point cloud registration (Steuernagel et al., 2023), Bayesian inference (Nguyen and Mueller, 2024), generative modelling (Atanackovic et al., 2025; Haviv et al., 2025), document analysis (Yurochkin et al., 2019), Gromov–Wasserstein approximations (Mémoli, 2011; Piening and Beinert, 2025a), and reinforcement learning (Ziesche and Rozo, 2023). Extending this framework with another Polish space $\mathcal{Y}$ for dataset labels yields the OT dataset distance (OTDD), defined on $\mathcal{P}_{2}(\mathcal{Y}\times\mathcal{P}_{2}(\mathcal{X}))$ (Alvarez-Melis and Fusi, 2020). However, all these approaches incur high computational cost due to repeated pairwise Wasserstein evaluations.
@@
Due to the complexity of OT, sliced Wasserstein distances (Bonneel et al., 2015; Nguyen, 2025) provide efficient OT-based alternatives to standard Wasserstein distances. Initially developed for probability measures on Euclidean spaces, they have since been extended to the sphere (Bonet et al., 2023a; Quellmalz et al., 2023), manifolds (Bonet et al., 2025b), functions (Garrett et al., 2024), hyperbolic spaces (Bonet et al., 2023b), the rotation group (Quellmalz et al., 2024), and matrices (Bonet et al., 2023c). Repeated sliced Wasserstein distances can be linearized by mapping measures onto function spaces (Naderializadeh et al., 2021; Nguyen, 2025). For WoW-type distances, sliced accelerations have been proposed for Gaussian mixtures (Nguyen and Mueller, 2024; Piening and Beinert, 2025b) and more generally for measures over measures (meta-measures) via sliced Wasserstein Busemann Gaussian (SWBDG/SWB1DG) distances (Bonet et al., 2025a) and the sliced OTDD (s-OTDD) (Nguyen et al., 2025). The SWBDG and SWB1DG, which appeared after the original submission of our work, are based on Busemann functions whose level sets act as a natural generalization of affine hyperplanes in the space of meta-measures. The numerical implementation relies on Gaussian approximations and a closed-form of the Busemann function. As an alternative, the s-OTDD employs a hierarchical slicing approach based on the method of moments. However, it is only well-defined for finite moments, and practical implementations are limited to the first few moments because of numerical instability – originally, the first five.
@@
  • Beyond 1d meta-measures, we extend our approach to the multivariate case by introducing the double-sliced Wasserstein (DSW [^1] Note that the acronym DSW is also used for the distributional sliced Wasserstein distance (Nguyen et al., 2021a), which modifies the slicing distribution of the classical sliced Wasserstein distance.) metric between meta-measures. Illustrating the usefulness of our DSW metric as a WoW replacement, we prove a form of topological metric equivalence between the two for discretized meta-measures.
@@ 2 Wasserstein Distances
The so-called Wasserstein distance or Kantorovich–Rubinstein metric is an optimal transport-based similarity gauge between probabilities on a common Polish space. To this end, let $\mathcal{X}$ be a Polish space, let $\mathcal{P}(\mathcal{X})$ be the space of Borel probability measures on $\mathcal{X}$ with respect to the Borel $\sigma$-algebra induced by the underlying metric, and let $\mathcal{P}_{p}(\mathcal{X})\coloneqq\{\mu\in\mathcal{P}(\mathcal{X})\mid\exists x_{1}\in\mathcal{X}:\int_{\mathcal{X}}d^{p}(x_{1},x_{2})\,\mathrm{d}\mu(x_{2})<\infty\}$, $p\in[1,\infty)$, be the subset of measures with finite $p$th moment. For $\mu\in\mathcal{P}(\mathcal{X})$ and a second Polish space $\mathcal{Y}$, the push-forward by a measurable map $T\colon\mathcal{X}\to\mathcal{Y}$ is defined by $T_{\sharp}\,\mu\coloneqq\mu\circ T^{{-}1}\in\mathcal{P}(\mathcal{Y})$. The set of transport plans between $\mu\in\mathcal{P}(\mathcal{X}),\nu\in\mathcal{P}(\mathcal{Y})$ is given by
@@
defines a metric on $\mathcal{P}_{2}(\mathcal{X})$. More precisely, $(\mathcal{P}_{2}(\mathcal{X}),\operatorname{W})$ is again a complete separable metric space, allowing the construction of the so-called Wasserstein over Wasserstein (WoW) distance ${\operatorname{\mathbf{W}}}(\cdot,\cdot;\mathcal{X})\coloneqq\operatorname{W}(\cdot,\cdot;\mathcal{P}_{2}(\mathcal{X}))$, which is studied in (Bonet et al., 2025c).
In difference to other similarity gauges like the Kullback–Leibler divergence or the total variation, Wasserstein distances leverage the underlying geometry, allowing for meaningful comparisons between empirical measures. Although the Wasserstein distance relies on a linear program, the actual calculation is computationally costly. For two empirical measures supported at $n$ points in $\mathbb{R}^{d}$ equipped with the Euclidean metric, the exact computation has complexity $\mathcal{O}(n^{3}\log n)$. The approximate computation based on entropic regularization still has complexity $\mathcal{O}(n^{2}\log n)$, see (Peyré and Cuturi, 2019). Notably, this computational burden becomes even more involved for non-Euclidean metric spaces, where the computation of the underlying distance itself is challenging. For instance, the computation of the WoW distance relies on the pointwise evaluations of Wasserstein distances. If the empirical meta-measures in $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ are supported on $N$ empirical measures on $\mathcal{P}_{2}(\mathbb{R}^{d})$, each with $n$ support points, then the approximate calculation of the required distance matrix already has complexity $\mathcal{O}(N^{2}n^{2}\log n)$.
@@ 3.1 Slicing Infinite Dimensional Banach Spaces
As preliminary step towards an implementable sliced WoW distance, we consider the slicing on an infinite dimensional, separable Banach space $U$ with norm $\lVert\cdot\rVert$. We denote the continuous dual by $U^{*}$ and equip it with the dual norm, i.e., $\lVert v\rVert\coloneqq\sup\{\lvert v(u)\rvert\mid u\in U,\lVert u\rVert\leq 1\}$ for all $v\in U^{*}$. Relying on the dual pairing, we generalize the slicing operator equation 5 by
$\pi_{v}\colon U\to\mathbb{R},\;u\mapsto\langle v,u\rangle{\coloneqq v(u)},\qquad v\in U^{*}.$ (7)
The crucial point in defining a sliced distance on $\mathcal{P}_{2}(U)\coloneqq\{\mu\in\mathcal{P}(U)\mid\int_{U}\lVert u\rVert^{2}\,\mathrm{d}\mu(x)<\infty\}$. is that there exists no uniform probability measure on the infinite dimensional sphere. As remedy, we choose an arbitrary $\xi\in\mathcal{P}_{2}(U^{*})\coloneqq\{\zeta\in\mathcal{P}(U^{*})\mid\int_{U^{*}}\lVert v\rVert^{2}\,\mathrm{d}\zeta(v)<\infty\}$ and define the $\xi$-based sliced Wasserstein distance as
@@
This approach extends the slicing on Hilbert spaces (Han, 2023). However, unlike (Han, 2023), we do not construct specific measures on the sphere. This allows the use of easy-to-sample slicing measures. If the support of $\xi$ covers all directions in $U^{*}$, the $\xi$-based SW distance becomes a metric. Here, the crucial point is the definiteness. For this, we show the Lipschitz continuity of the integrant in equation 8 and exploit the uniqueness of the characteristic function, see Appendix A.
@@
For $\xi\in\mathcal{P}_{2}(U^{*})$, the $\xi$-based SW distance defines a well-defined pseudo-metric. If $\operatorname{supp}\xi\cap\operatorname{span}v\not\in\bigl\{\emptyset,\{0\}\bigr\}$ for all $v\in U^{*}\setminus\{0\}$, then equation 8 defines a metric on $\mathcal{P}_{2}(U)$.
If the slicing measure $\xi$ has full support, then the assumption in Theorem 3.1 is fulfilled, and the $\xi$-based SW distance is a metric. Two measures $\xi_{1},\xi_{2}\in\mathcal{P}_{2}(U^{*})$ are equivalent if they are mutually absolutely continuous. If their Radon–Nikodým derivatives ${\,\mathrm{d}\xi_{1}}/{\,\mathrm{d}\xi_{2}}$ and ${\,\mathrm{d}\xi_{2}}/{\,\mathrm{d}\xi_{1}}$ are bounded, then the resulting SW distances are metrically equivalent. The proof is given in Appendix A.
@@
In the finite-dimensional Euclidean setting, special cases of the $\xi$-based SW distance correspond, for instance, to so-called energy measures on the sphere (Nguyen and Ho, 2023) and the standard Gaussian (Nadjahi et al., 2021). Relying on the latter, we obtain a strong equivalence to original SW equation 6 if $\xi$ is equivalent to the standard Gaussian. The short proof, which relies on Proposition 3.2 and the fact that the classical SW distance equation 6 and the reference-based SW distance equation 8 with the standard Gaussian reference coincide on $\mathcal{P}_{2}(\mathbb{R}^{d})$, is given in Appendix A.
@@
$c_{1}\operatorname{SW}(\mu,\nu;\xi)\leq{\operatorname{SW}(\mu,\nu;\eta)}=\operatorname{SW}(\mu,\nu)\leq c_{2}\operatorname{SW}(\mu,\nu;\xi)\quad\forall\mu,\nu\in\mathcal{P}_{2}(\mathbb{R}^{d}).$ (10)
@@ 3.2 Slicing the 1D Wasserstein Space
For the later implementation, we require an easy-to-sample slicing measure. To this end, we propose to use Gaussians, i.e., measures $\xi\in\mathcal{P}_{2}(L_{2}([0,1]))$ such that $\pi_{v,\sharp}\xi$ is Gaussian for all $v\in L_{2}([0,1])$, see (Bogachev, 1998). On $L_{2}([0,1])$, there exists a one-to-one correspondence between Gaussian measures and Gaussian processes (Rajput and Cambanis, 1972, Thm. 2). In our numerics, we restrict ourselves to the Gaussian process $G$ that is related to the covariance kernel
@@
Since the kernel is smooth, the sample paths (realizations) of $G$ are smooth too, i.e., $G~\in~\mathcal{C}^{\infty}([0,1])$ almost surely, see (Da Costa et al., 2026, Cor. 1). Since $k_{\sigma}$ is universal (Steinwart, 2001), the corresponding Gaussian measure has full support (Van Der Vaart et al., 2008).
@@ 4 Double-Slicing the Wasserstein Space
where we again integrate with respect to the uniform measure. Based on finite-dimensional slicing, $\operatorname{\mathbf{SW}}$} essentially reduces meta-measures $\bm{\mu}$ and $\bm{\nu}$ to a series of 1d meta-measures $\bm{\pi}_{\theta,\sharp}\bm{\mu}$ and $\bm{\pi}_{\theta,\sharp}\bm{\nu}$.
As the computation of the 1d WoW distance remains challenging, we resort to a hierarchical slicing approach, inspired by (Nguyen et al., 2025; Piening and Beinert, 2025b). Using quantile slicing equation 11 with slicing measure $\xi\in\mathcal{P}_{2}(L_{2}([0,1]))$, we introduce the double-sliced WoW distance
@@
The detailed proofs are given in Appendix B. There, a transport plan is constructed to rely on the metric properties of WoW. The equivalence is then proven via a discretization of the Gaussian process and the compactness of $\mathcal{P}_{\mathrm{e}}^{N}(\mathcal{P}_{\mathrm{e}}^{\tilde{n}}(\mathcal{X}))$. Numerically, $\operatorname{\mathbf{DSW}}$} can be implemented combining several integration techniques, see Appendix C.1 and C.2. Here, we consider the Gaussian $\xi$ related to the kernel in equation 12 and empirical meta-measures $\bm{\mu},\bm{\nu}\in\mathcal{P}_{\mathrm{e}}(\mathcal{P}_{\mathrm{e}}(\mathbb{R}^{d}))$ as in equation 17. To approximate the outer integral over $\mathbb{S}^{d-1}$ and the inner integral over $\xi$ simultaneously, we employ the Monte Carlo method. For a sample path $g$ of the corresponding Gaussian process and a direction $\theta\in\mathbb{S}^{d-1}$, we evaluate the integrand as follows: The first slicing gives $\bm{\pi}_{\theta,\sharp}\bm{\mu}=(1/N)\sum{}_{i=1}^{N}\delta_{\pi_{\theta,\sharp}\mu_{i}}$, and the quantile mapping $q_{\sharp}\bm{\pi}_{\theta,\sharp}\bm{\mu}=(1/N)\sum{}_{i=1}^{N}\delta_{q(\pi_{\theta,\sharp}\mu_{i})}$, where the piecewise constant quantile functions can be determined by sorting the support points. For the slicing on $L_{2}([0,1])$, we employ a quadrature to approximate the inner product. To this end, for knots $t_{1},\dots,t_{R}\in[0,1]$ and weights $w_{1},\dots,w_{R}$, we estimate
$\widehat{\pi_{g,\sharp}q_{\sharp}\bm{\pi}_{\theta,\sharp}\bm{\mu}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\widehat{\langle q(\pi_{\theta,\sharp}\mu_{i}),g\rangle}}\quad\text{with}\quad\widehat{\langle q(\pi_{\theta,\sharp}\bm{\mu}_{i}),g\rangle}=\sum_{r=1}^{R}w_{r}\,q(\pi_{\theta,\sharp}\mu_{i})(t_{r})\,g(t_{r}).$ (19)
@@
$\widehat{{\operatorname{\mathbf{DSW}}}}(\bm{\mu},\bm{\nu})\coloneqq\biggl({\frac{1}{S}\sum_{s=1}^{S}\operatorname{W}^{2}\bigl(\widehat{\pi_{g_{s},\sharp}q_{\sharp}\bm{\pi}_{\theta_{s},\sharp}\bm{\mu}},\widehat{\pi_{g_{s},\sharp}q_{\sharp}\bm{\pi}_{\theta_{s},\sharp}\bm{\nu}};\mathbb{R}\bigr)}\biggr)^{\frac{1}{2}}.$ (20)
@@
At its core, $\operatorname{\mathbf{DSW}}$} requires the computation of the domain projection $\bm{\pi}_{\theta}$ and a quantile projection $q$. As an alternative approach, s-OTDD considers a variation of $\bm{\pi}_{\theta}$ with general 1d feature projections, including convolutions, and a (finite) moment projection instead of quantiles (Nguyen et al., 2025). Another concurrent work constructs a direct slicing of $\mathcal{P}_{2}(\mathcal{P}_{2}(\mathbb{R}^{d}))$ via so-called Busemann projections onto geodesic rays and relies on Gaussian approximations (Bonet et al., 2025a). Lastly, notice that the collection of these push-forward $q_{\sharp}\bm{\pi}_{\theta,\sharp}$ can be interpreted as meta-measure version of the so-called sliced-Wasserstein embedding (Kolouri et al., 2016; Naderializadeh et al., 2021; Nguyen, 2025) that maps a classical measure in $\mathcal{P}_{2}(\mathbb{R}^{d})$ to the quantile functions of its slices.